OpenStudy (anonymous):

A bag contains 10 beads --2 black, 3 white, and 5 red. A bead is selected at random. Find the probability of selecting 3 non red beads without replacement. Determine whether the events are independent or dependent.

OpenStudy (anonymous):

first non red prob is \(\frac{5}{10}\) second no red given first non red, you have 4 not red, 9 in the bag so \(\frac{4}{9}\) this non red given first two you have 3 not red, and 8 in the bag so \(\frac{3}{8}\) now multiply all these together

OpenStudy (anonymous):

i am not exactly sure how the answer the second question but it think you are supposed to say "dependent" because you do not replace the beads, so second draw depends on what happens the first draw

OpenStudy (anonymous):

i got 1/12

OpenStudy (anonymous):

me too. and yes, it is dependent.